# CriticEval: Evaluate Large Language Model as Critic

### v4

@@ F.2 Correct or Golden Response Generation
Appendix G Human Annotation Details
G.1 Evaluation Protocol
Three to five human annotators annotate each scalar-valued critique in each task. These human annotators first annotate each textual critique, which are summarized by another supervisor annotator. To reduce the bias of these human annotators, we designed a rigorous data annotation verification process to iteratively train these annotators, thereby ensuring the stability and reliability of the annotation quality. Biases among human annotators may arise from factors such as the annotator’s gender and professional background. To minimize the impact of annotators’ biases on the quality of our human annotations, we first selected a diverse group of annotators from a crowd-sourcing platform to annotate each critique sample collaboratively.
All raters have been paid adequate wages. The hourly wage of our human annotators is about 5.69 USD, which is much higher than average hourly wage 3.13 USD on Amazon Mechanical Turk (hara2018data, ).
G.2 Score Rubrics for Different Tasks
@@ G.3 Internet Search
G.4 Exclude Mention of Ground-Truth
@@
G.5 Statistics of Human Annotation on Reference Critiques
@@
Appendix H Case Study
@@ H.1 Examples in CriticEval
H.2 Case Study Proves Reference Critique
@@
H.3 Subjective Evaluation Cases
@@
H.4 Objective Evaluation Cases
@@
Appendix I Analysis about Length Bias in Subjective Evaluation
Appendix J Overall Score Computation
Appendix K Complete Results
K.1 Complete Results for Different Qualities
@@ Appendix I Analysis about Length Bias in Subjective Evaluation
K.2 Complete Results for Different Tasks
@@
Appendix L Evaluated LLMs
@@
L.1 Instruction-tuned LLMs
@@
L.2 Critique-tuned LLMs
@@
Appendix M Interpretable Analysis of the Quality of Textual Critiques
@@
Appendix N Error Patterns in Responses
@@
Appendix O Likert Score for Responses
@@
Appendix P Visualization of Relationship between Model Scales and Critique Ability
@@ 18. Answer: [Yes]
  19. Justification: Our experimental results could be easily reproduced because we have already discussed the necessary details in Section 6.1, Appendix H and Appendix K.
@@ 28. Answer: [Yes]
  29. Justification: We have well described all the details about inference procedure in Section 6, Appendix G, Appendix F, Appendix E, Appendix D, Appendix L and Appendix C.
@@ 38. Answer: [Yes]
  39. Justification: The time cost, compute worker, CUDA memory for LLM inference in our proposed CriticEval dataset are well described in Appendix L.
@@ 68. Answer: [Yes]
  69. Justification: We have discussed this part of the content carefully in Appendix G.
@@ 73. Answer: [Yes]
  74. Justification: The participants in our work are asked to evaluate the quality of text generations, which doesn’t raise any potential risks. Therefore, the Institutional Review Board (IRB) approvals are easily obtained for our work.

### v5

@@ F.2 Correct or Golden Response Generation
Appendix G Reasons of Utilizing Human-in-the-Loop Data Generation Pipeline
We construct the high-quality reference critiques by using human-in-the-loop data generation pipeline, which is motivated by two essential considerations (effectiveness and efficiency).
Human-written Critiques are Usually Insufficient - Sub-optimal (Effectiveness)
Our trial human annotation reveals that human annotators might neglect some apparent or severe issues when writing critiques from scratch, consistent with findings in recent studies from OpenAI mcaleese2024llmcriticshelpcatch ; saunders2022selfcritiquing . Our experimental result in Appendix M demonstrate that neglecting issues usually leads to low-quality critiques. In contrast, despite the possibility of generating wrong critiques, LLMs like GPT-4 offer more comprehensive and detailed critiques mcaleese2024llmcriticshelpcatch . By revising LLM’s errors by human experts, the final critiques could be more comprehensive and accurate, leveraging the strengths of both human annotators and LLMs.
Annotating Challenging Critique Task from Scratch Cost A Lot (Efficiency)
Writting critiques from scratch is a significant challenge liu2023alignbench ; wang2023shepherd . For instance, Shepherd fengshenbang incurred an annotation cost of 8\$ per sample, leading to over 28,864\$ and 1,350 work hours to annotate the entire CriticEval, which is unbearable for our project. Thus, we have to employ advanced LLMs to generate draft critiques, followed by human annotation. GPT-4 was chosen because our preliminary studies indicate that it is the most reliable LLM for producing draft critiques, while other LLMs are much worse. Consequently, a diverse set of LLMs introduces more noise in generated critiques, bringing more difficulties to human annotators.
In conclusion, the human-in-the-loop pipeline achieves the trade-off between annotation cost and quality. We promise to add these details to the revised paper to emphasize the motivation of using the human-in-the-loop pipeline.
Appendix H Human Annotation Details
H.1 Evaluation Protocol
Three to five human annotators annotate each scalar-valued critique in each task. Biases among human annotators may arise from factors such as the annotator’s gender and professional background. To minimize the impact of annotators’ biases on the quality of our human annotations, we first selected a diverse group of annotators from a crowd-sourcing platform to annotate each critique sample collaboratively. All human annotators have been paid adequate wages. The hourly wage of our human annotators is about 5.69 USD, which is much higher than average hourly wage 3.13 USD on Amazon Mechanical Turk (hara2018data, ).
Before annotation, we designed a rigorous data annotation verification process to iteratively train these annotators, thereby ensuring the stability and reliability of the annotation quality. During annotation, these human annotators first annotate each textual critique, which are summarized by another supervisor annotator. After annotation, the supervisors (authors in our paper) conducted a 5% sample inspection. If the error rate exceeds the threshold, annotators are asked to revise their work until the error rate is lower than the threshold.
H.2 Inspect Errors in Datasets
It should be noted that some underlying datasets contain inaccuracies that may lead to compounding effects during evaluation. For example, we have noticed some incorrect solutions and rationales for mathematics and coding questions during our human evaluation process. In our work, to mitigate the effects of such errors, human annotators are asked to meticulously examine each question, the provided golden answers (only for mathematics and coding tasks) and the evaluated responses and critiques. They are asked to exclude instances where the golden answers or questions are flawed or incorrect, like wrong solutions to mathematics and coding questions.
H.3 Score Rubrics for Different Tasks
@@ H.4 Internet Search
H.5 Exclude Mention of Ground-Truth
@@
H.6 Statistics of Human Annotation on Reference Critiques
@@
Appendix I Case Study
@@ I.1 Examples in CriticEval
I.2 Case Study Proves Reference Critique
@@
I.3 Subjective Evaluation Cases
@@
I.4 Objective Evaluation Cases
@@
Appendix J Analysis about Length Bias in Subjective Evaluation
Appendix K Overall Score Computation
@@
Appendix L Complete Results
L.1 Complete Results for Different Qualities
@@
L.2 Complete Results for Different Tasks
@@
Appendix M Human Performance in CriticEval
In this section, we provide more details and comparison between LLMs and human performance. Specifically, we conduct the human annotation of the subjective tasks on the CriticEval test set, and the overall human-level performance is shown in Table 29. Note that the cohort and corresponding set of human critiques does not represent the best possible human performance; instead, they represent the capability of annotators selected for this human performance annotation of the CriticEval test set.
Table 29: Comparison between Human Performance and GPT-4-turbo.
$\boldsymbol{F_{s}}$ Sub. $\boldsymbol{F_{s}}$ Obj. $\boldsymbol{F_{c}}$ Sub. $\boldsymbol{F_{c}}$ Obj. $\boldsymbol{CR}$ Sub. $\boldsymbol{CR}$ Obj.
GPT-4 7.84 63.54 7.89 57.33 7.69 69.67
Human 5.61 67.69 5.22 60.67 6.63 75.69
It can be found that the human-level significantly outperforms GPT-4 on the objective task, while it is inferior to GPT-4 on the subjective evaluation. Therefore, we conduct the Quantitative and Qualitative Analysis to understand the performance gap between humans and GPT-4 in subjective evaluation.
Quantitative Analysis
We conduct the fine-grained failure modes analysis, and the distribution of each failure mode for feedback, comparison and correction dimensions are shown in Table 30, Table 31 and Table 32. The numbers in following tables indicate the frequencies of error types in GPT-4 and human-written critiques. The detailed description of each failure mode can be found in Section 6.8. As for the feedback and comparison dimensions, the distribution of E1 (missing issues), E2 (missing suggestions or low-quality suggestions), and E7 (insufficient analysis) in human-written critiques is significantly higher than that of GPT-4. In contrast, the distribution of other error types is significantly lower. As for the correction dimension, human annotators usually do not follow suggestions in the provided feedback (E9) and generate additional errors (E11). Through communicating with the annotators, we notice that the primary cause of this issue is that some tasks require domain-specific knowledge, and the lack of this knowledge among human annotators results in lower-quality corrections. In summary, the human-written critiques are often less comprehensive than GPT-4, significantly reducing the quality. In contrast, the mistakes in human-written critiques are significantly less than that of GPT-4. This phenomenon is consistent with our preliminary study and recent findings mcaleese2024llmcriticshelpcatch ; saunders2022selfcritiquing , further proving the effectiveness and reasonableness of leveraging the human-in-the-loop pipeline to construct comprehensive and accurate reference critiques.
Table 30: Comparison between Human Performance and GPT-4-turbo in feedback dimension ($\boldsymbol{F_{s}}$).
E1 E2 E3 E4 E5 E6 Other
GPT-4 17.99 18.71 16.37 15.83 10.07 14.93 6.12
Human 21.18 24.48 11.36 15.27 9.06 12.13 6.52
Table 31: Comparison between Human Performance and GPT-4-turbo in comparison dimension ($\boldsymbol{F_{c}}$).
E1 E2 E3 E4 E5 E6 E7 E8 Other
GPT-4 16.67 11.02 11.29 15.59 4.30 6.99 19.35 12.10 2.69
Human 19.71 15.29 7.65 9.51 3.82 4.80 24.90 11.67 2.65
Table 32: Comparison between Human Performance and GPT-4-turbo in correction dimension ($\boldsymbol{CR}$).
E1 E2 E3 Other
GPT-4 23.46 43.83 21.60 11.11
Human 29.38 36.88 25.00 8.75
Qualitative Analysis
We inspect the human-written critiques in the subjective evaluation tasks to understand the source of the performance difference. In general, human annotators write fewer comments than LLMs, and the comments are usually general and brief. Besides, many tasks involve domain-specific knowledge that humans may lack, but GPT-4 excels in (albeit with potential hallucinations).
Appendix N Evaluated LLMs
@@
N.1 Instruction-tuned LLMs
@@
N.2 Critique-tuned LLMs
@@
Appendix O Interpretable Analysis of the Quality of Textual Critiques
@@
Appendix P Error Patterns in Responses
@@
Appendix Q Likert Score for Responses
@@
Appendix R Visualization of Relationship between Model Scales and Critique Ability
@@
Appendix S How Few-shot Examples Affect Performance
In this paper, we have studied the few-shot prompting strategy. However, our results demonstrate that few-shot prompting reduces performance across various LLMs. As illustrated in Table 34, when using 1-5 examples in objective feedback evaluations, we observed a significant decline in LLMs’ performance as the number of examples increased. This intriguing phenomenon may be due to the complexity of the critique task, where few-shot examples might impede the LLM’s understanding of the evaluated responses. Consequently, CriticEval currently does not utilize few-shot prompting by default. Given the emerging interest in critique ability research, we look forward to future works investigating advanced inference strategies to improve critique ability of LLMs.
Table 34: The critique performance (Spearman correlation) of LLMs by few-shot prompting.
Models No. Few-shot 1 2 3 4 5
Llama-3-7B-Chat 61.34 58.13 54.25 52.99 53.23 50.11
InternLM2-20B-Chat 69.86 66.26 64.99 63.32 60.33 61.72
Appendix T Scalability and Cost about CriticEval
In this section, we provide the cost of constructing one task and inferencing one LLM in CriticEval.
T.1 Construction Cost
Collect Evaluated Responses for All Tasks
  • Open-source LLMs: a GPU server with 8 A100 (80G) cards is used to generate evaluated responses, and the total GPU hours are 4.26 hours, approximately 82.88\$ (refer to the price of Alibaba Cloud).
  • Closed-source LLMs: the average cost for each LLM is 0.89\$.
Generate and Revise GPT-4 Critiques
The cost of the human annotation is computed under these settings: (1) Four human annotators (3 annotators and one supervisor); (2) 5.69\$ hourly wage for each annotator; (3) Average 400 samples in one task. The overall construction cost are shown in Table 35, which is affordable dubois2024alpacafarm .
Table 35: The cost of constructing the critiques.
For Each New Task Cost (\$) Time (hour)
Generate Critiques (GPT-4) 3.09 -
Human Annotation 303.53 53.34
Overall 306.62 53.34
T.2 Average Computation Cost for One LLM
As shown in Table 36The overall cost of the test and dev set is 13.19+9.94=23.13\$, comparable to the evaluation cost on the AlpacaEval benchmark (5-15\$) alpaca_eval . These costs are essential for CriticEval, as they guarantee the reliability of critique evaluation. We promise to add these details to the Appendix of our revised submission.
Table 36: The computation cost of inference one LLM in CriticEval.
Dimensions Cost of Test set (\$) Cost of Dev set (\$)
Feedback 4.21 5.09
Correction 2.11 2.67
Comparison 3.62 5.43
Overall 9.94 13.19
Appendix U Multilingual Support
The primary goal of CriticEval in the current stage is to construct a reliable and comprehensive evaluation for critique ability. We agree that it is essential to study multilingual critiques and intend to broaden CriticEval to include other languages in future work. The following content briefly introduces our preliminary solution on how to achieve this goal.
Construct Multilingual CriticEval
Following the previous work shi2023language , CriticEval could be translated to various languages, especially low-resource languages, with human annotation for revising translation inaccuracies. the most direct way is to translate CriticEval into various languages, with human annotation for revising translation inaccuracies.
Evaluate Multilingual CriticEval
While the reliability of objective evaluation could be ensured, the reliability of subjective evaluation is limited by the multilingual capability of the judge model (GPT-4). We recommend back-translating multilingual critiques into English and evaluating them within English CriticEval.
@@ 18. Answer: [Yes]
  19. Justification: Our experimental results could be easily reproduced because we have already discussed the necessary details in Section 6.1, Appendix I and Appendix L.
@@ 28. Answer: [Yes]
  29. Justification: We have well described all the details about inference procedure in Section 6, Appendix H, Appendix F, Appendix E, Appendix D, Appendix N and Appendix C.
@@ 38. Answer: [Yes]
  39. Justification: The time cost, compute worker, CUDA memory for LLM inference in our proposed CriticEval dataset are well described in Appendix N.
@@ 68. Answer: [Yes]
  69. Justification: We have discussed this part of the content carefully in Appendix H.
@@ 73. Answer: [Yes]
  74. Justification: We have submitted an application to the Institutional Review Board (IRB). After evaluation, the committee determined that the human annotation for critique task posed no ethical concerns, and accordingly, we conduct the human annotation. Besides, we have adequately addressed minimal ethical concerns by following efforts: ensuring the fairness of the data collection by selecting a diverse group of annotators, providing clear guidelines, and compensating the annotators well, as described in Appendix B and Appendix H. Therefore, these efforts make sure our work does not violate the NeurIPS Code of Ethics.